bdrammeh:

a 2.99 x 10^-6 C charge is moving perpendicular (90) to the earth magnetic field (5.00 x 10^-5 T). If the force on it is 2.14 x 10^-8 N, how far is it moving ? m/s

TheSmartOne:

Hi! So do you know the formula we should be using?

bdrammeh:

no

TheSmartOne:

F = qvB sin\(\theta\) Have you seen this formula before? q is the charge v is the velocity B is the magnetic field \(\theta\) is the angle between the velocity and magnetic field

bdrammeh:

YES

bdrammeh:

PLUG IT IN FOR ME

bdrammeh:

AND THEN ILL ANSWER IT

TheSmartOne:

okay okay I'll do it for you this one time, but there's a reason why I haven't been Plugging in is the easy part and you should be able to do it yourself. Figuring out the equation requires practice and all But if you can't plug it in, or figure out what to plug in, then it'll be challenging as you continue your physics, chemistry, math classes But here we go The question says the charge = q = 2.99*10^-6 C Magnetic field = B = 5.00 * 10^-5 T \(\theta = 90^o\) And we're solving for velocity or v Do you follow so far?

bdrammeh:

I GUESS

bdrammeh:

just answer this one for me and ill figure the next one

TheSmartOne:

So plug and chug Forgot the force but it's also given in the question which is 2.14*10^-8 N F = qvB sin\(\theta\) 2.14*10^-8 = (2.99*10^-6) * v * (5.00*10^-5) sin\((90^o)\) So solve for v. Hint: sin 90 = 1

TheSmartOne:

Let me know what you get First multiply the numbers on the right side of the equal sign And then divide by that number on both sides so that way you can get 'v' all by itself

bdrammeh:

143

TheSmartOne:

That's right!